OpenStudy (anonymous):

Create a word problem of your own that involves fractions or mixed numbers. Write a description of how to use the problem-solving strategy Work Backward to solve it. 4 = 1 + 1 + ½+ ½+ 1 = (½+½) + (½+½) + 1 + 1=4 Is this ^ correct for the question above?

OpenStudy (anonymous):

Your equation is good but you are missing the word problem part. A word problem is a story that gives the information. Example: You and four of your friends are going to the movies. You buy 5 bags of popcorn, one for each of you. You and two of your friends eat all of your popcorn while your other two friends only ate half of their bags. How many full bags of popcorn were eaten? Now you must separate the information. # of bags of popcorn in all = 5 # of bags of popcorn eaten = 4 # of people = 5 5 people = 1 + 1 + 1 + 1 + 1 Only 4 bags eaten = 5 - 1 = 4 4 = 1+ 1 + 1 + 1 but there are 5 people in all, so you must have 5 values. 4 = 1 + 1 + 1 + 1/2 + 1/2 4 = 1 + 1 + 1 + (1/2 + 1/2) I hope this helps.